-6(3-3a)-8(6a+5)=32
Answer by addingup(3677)

-6(3-3a)-8(6a+5) = 32
18a-18-8(6a+5) = 32
18a+-40-48a-18 = 32
Group like terms
(18a-48a)+(-18-40) = 32
Simplify
-30a+(-18-40) = 32
-58-30a = 32
Add 58 on both sides
-30a = 90
Divide both sides by -30
a = -3
:
John